how do you solve for ex^2

Tactor 20 T 180 SS 35C n= 26 Zx2- Non cuntact SSこ350
Answer #1

\sum X_i^2 = (n-1)SS + n \overline{X}^2

=19*380 + 20*9^2

= 8840

19*390 + 20*4^2

= 7730

19*350 + 20*9^2

= 8270

19*400 + 20*8^2

= 8880

hence

\sum X^2

= 8840 + 7730+ 8270 + 8880

= 33720